OpenStudy (anonymous):

Can some explain how to do this:

OpenStudy (anonymous):

Write the equation of a line that is parallel to 9x + 3y = 15 and goes through the point (–1, 2)

OpenStudy (anonymous):

first you need the slope of \(9x+3y=15\) so you can use the point slope formula do you know what that slope is?

OpenStudy (anonymous):

point slope= y=mx+b

OpenStudy (anonymous):

yes, and you need \(m\) to you know the slope of \(9x+3y=15\) ? you need that number

OpenStudy (anonymous):

how thogh?

OpenStudy (anonymous):

you can do it by solving for \(y\): \[9x+3y=15\] \[3y=-9x+15\] \[y=-3x+5\] so the slope is \(-3\)

OpenStudy (anonymous):

I think I know it can I show u guys

OpenStudy (anonymous):

satellite73 I know what r u doing but she have to do focus in class more 3y=-9x+15 write it y=mx+b and than look for b2 of the second equation y2 which have the same m using this information. y(-1)=2

OpenStudy (anonymous):

or you can remember (memorize) that the slope of \(ax+by=c\) is \(-\frac{a}{b}\)

OpenStudy (anonymous):